Unit Labor Require ments Widge ts Cloth Home 100 200 Foreign 60 30

1) Given the information in the table above. What is the opportunity cost of Cloth in terms of Widgets in Home?

2) Given the information in the table above. If these two countries trade these two goods in the context of the Ricardian model of comparative advantage, then what is the lower limit of the world equilibrium price of widgets?

3) Given the information in the table above. If these two countries trade these two goods with each other in context of the Ricardian model of comparative advantage, what is the lower limit for the price of cloth?


4) Given the information in the table above. What is the opportunity cost of cloth in terms of Widgets in Home?
Answer #1

Q1) for home

In order to produce 100 units of cloth, give up 200 widgets

To produce 1 cloth, give up 200/100 = 2 W.

So Opportunity cost of 1 C = 2W

Q2) now country with the lower opportunity cost of production of cloth in terms of W, will have CA in cloth Production.

So for foreign, Opportunity cost of cloth = 30/60 = .5 W

Thus foreign has lower Opportunity cost of production of cloth ,

so foreign has CA in cloth

thus home has CA in widgets

now opportunity cost of W in home = .5C

In foreign, OC = 2C

So lower limit of W in terms of cloth = .5 cloth

3) lower limit of C in terms of W = .5 units of widgets

4) same answer as in part 1)

So Opportunity cost of C in terms of W, at home = 2 units of widgets